OpenStudy (anonymous):

help me please! * draw a line with the given slope, 'm' , through the point given. 19. (-4,5) ; m= 2/3 20. (3,1) ; m=0 21. (-3,-4) ; m= -2

OpenStudy (anonymous):

y=mx+b just plug in the numbers to solve for b

OpenStudy (anonymous):

give me ex :)

OpenStudy (anonymous):

19. 5=2/3 (-4)+b 20. 1=0(3)+b 21. -4= (-2)(-3)+b

OpenStudy (anonymous):

solve for b and then rewrite as y=mx+b with m and b filled in

OpenStudy (anonymous):

how do i draw line from that?

OpenStudy (anonymous):

with the slope m=2/3 u can go 2 up 3 right so (3,2) (-4,5) for 19

OpenStudy (anonymous):

once you find b, then you have everything you need. if you just need to draw it, you have the slope and a point.

OpenStudy (anonymous):

1=0(3)+b = 1=0+b ?

OpenStudy (anonymous):

19 another point would be: -1, 7

OpenStudy (anonymous):

can you write me all of the points? please;)

OpenStudy (anonymous):

yes that is correct for 20

OpenStudy (anonymous):

so for 20 . = 1=b ?

OpenStudy (anonymous):

yes. so when x=0, y=1. so theres another point to help draw the line.

OpenStudy (anonymous):

what about 21. ?

OpenStudy (anonymous):

god this is so hard!

OpenStudy (anonymous):

-4=6+b

OpenStudy (anonymous):

can you write me all of the points? from 19-21 ?

OpenStudy (anonymous):

all of the intercepts? once u find b, then you have an intercept. (0,b)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you can do it. dont try to look to the end, think of what you need to get there asnd focus on those

OpenStudy (anonymous):

all you need to draw the line is 2 points

OpenStudy (anonymous):

what do you mean 2 points?

OpenStudy (anonymous):

i still dont understand anuthing!

OpenStudy (anonymous):

can you connect the dots? if yes then you can draw a line from 2 points

OpenStudy (anonymous):

i know nut i dont know the points...

OpenStudy (anonymous):